### v1

@@ 3 Proximal Operator
(26) $\operatorname{prox}_{\lambda r_{\gamma}}(\bm{y})=\arg\min_{\bm{x}\in\mathbb{R}^{n}}\gamma\bigl(\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}\bigr)+(1-\gamma)\|\bm{x}\|_{2}^{2}+\frac{1}{2\lambda}\|\bm{x}-\bm{y}\|_{2}^{2},$
@@
Theorem 3.1 (Solution’s Existence).
@@
Proof 3.2.
Clearly, the objective function in (26) is proper and closed. It follows from the inequality $\|\bm{x}\|_{1}\geq\|\bm{x}\|_{2}$ that $f_{\gamma,\alpha}(\bm{x})=\gamma\|\bm{x}\|_{1}-\gamma\alpha\|\bm{x}\|_{2}+(1-\gamma)\|\bm{x}\|_{2}^{2}+\frac{1}{2\lambda}\|\bm{x}-\bm{y}\|_{2}^{2}\geq\left(1-\gamma+\frac{1}{2\lambda}\right)\|\bm{x}\|_{2}^{2}+\left[\gamma(1-\alpha)-\frac{1}{\lambda}\|\bm{y}\|_{2}\right]\|\bm{x}\|_{2}+\frac{1}{2\lambda}\|\bm{y}\|_{2}^{2}$. The right-hand side is a quadratic function of $\|\bm{x}\|_{2}$ with a positive leading coefficient $1-\gamma+\frac{1}{2\lambda}>0$. Hence, $f_{\gamma,\alpha}(\bm{x})\to+\infty$ as $\|\bm{x}\|_{2}\to\infty$, which implies that $f_{\gamma,\alpha}$ is coercive. Since $f_{\gamma,\alpha}$ is also lower semicontinuous, its sublevel sets are compact and so $\operatorname{prox}_{\lambda r_{\gamma}}(\bm{y})\neq\emptyset$.
Lemma 3.3.
@@
    If $\|\bm{y}\|_{\infty}>\gamma\lambda$, then $\bm{x}^{*}=\frac{\|\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})\|_{2}+\alpha\gamma\lambda}{\bigl(1+2\lambda(1-\gamma)\bigr)\,\|\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})\|_{2}}\;\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})$, where $\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})=\operatorname{sign}(\bm{y})\odot\max\big(|\bm{y}|-\gamma\lambda,\,\bm{0}\big)$.
@@
    If $\|\bm{y}\|_{\infty}=\gamma\lambda$, then the optimal solution $\bm{x}^{*}$ satisfies $\bm{x}^{*}\in\Bigl\{\bm{x}\in\mathbb{R}^{n}:x_{i}=0\text{ for }|y_{i}|<\gamma\lambda,\ \|\bm{x}\|_{2}=\alpha\gamma\lambda,\ x_{i}y_{i}\geq 0,\forall i\Bigr\}$, and $\bm{x}^{*}$ is non-unique if multiple $|y_{i}|=\gamma\lambda$.
@@
    If $(1-\alpha)\gamma\lambda<\|\bm{y}\|_{\infty}<\gamma\lambda$, then $\bm{x}^{*}$ is $1$-sparse, supported on an index $i$ with $|y_{i}|=\|\bm{y}\|_{\infty}$, and $\bm{x}^{*}\in\Bigl\{\bm{x}\in\mathbb{R}^{n}\Big|\|\bm{x}\|_{2}=\|\bm{y}\|_{\infty}+(\alpha-1)\gamma\lambda,\ x_{i}y_{i}\geq 0\Bigr\}.$
@@
Proof 3.4.
By combining the quadratic terms in the proximal mapping (26), the resulting operator can be expressed in the same form as the $\ell_{1}-\alpha\ell_{2}$ proximal mapping $\operatorname{prox}_{\lambda r_{\gamma}}(\bm{\tilde{y}})=\arg\min_{\bm{x}\in\mathbb{R}^{n}}\,\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}+\frac{1}{2\tilde{\lambda}}\|\bm{x}-\bm{\tilde{y}}\|_{2}^{2}$, where $\tilde{\lambda}=\frac{\lambda\gamma}{1+2\beta(1-\gamma)}$ and $\bm{\tilde{y}}=\frac{1}{1+2\beta(1-\gamma)}\bm{y}$. This reformulation allows one to directly apply the closed-form expressions derived for the $\ell_{1}-\alpha\ell_{2}$ case. The remaining proof follows that of [lou2018fast] and is therefore omitted.
Remark 3.5.
When $\alpha=0$ and $\gamma=1$, the analytical solution of the proximal operator $\mathrm{prox}_{\lambda r_{\gamma}}$ degenerates to the soft-thresholding function $\mathcal{\bm{S}}_{\lambda}(\bm{y},\lambda)$ associated with the $\ell_{1}$ norm. When $\alpha=0$ and $\gamma\in(0,1)$, the analytical solution of $\mathrm{prox}_{\lambda r_{\gamma}}$ specializes to the soft-thresholding function of Elastic Net. When $\alpha>0$ and $\gamma=1$, the closed-form of $\mathrm{prox}_{\lambda r_{\gamma}}$ coincides with that of the $\ell_{1}-\alpha\ell_{2}$ proximal operator. This lemma can thus be viewed as a direct generalization of Lemma 1 in [lou2018fast] and Lemma 1 in [liu2017further].
Lemma 3.6.
@@
(27) $\bm{p}=\frac{1}{\lambda\gamma}\bm{y}-\left(\frac{1}{\lambda\gamma}+\frac{2(1-\gamma)}{\gamma}-\frac{\gamma\alpha}{\|\bm{x}^{*}\|_{2}}\right)\bm{x}^{*}.$
@@
By the subgradient inequality for $\|\cdot\|_{1}$, one has $\|\bm{x}\|_{1}\geq\|\bm{x}^{*}\|_{1}+\langle\bm{p},\bm{x}-\bm{x}^{*}\rangle$, which implies $\|\bm{x}^{*}\|_{1}-\|\bm{x}\|_{1}\leq\langle\bm{p},\bm{x}^{*}-\bm{x}\rangle$ and so $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq\gamma\langle\bm{p},\bm{x}^{*}-\bm{x}\rangle-\gamma\alpha\left(\|\bm{x}^{*}\|_{2}-\|\bm{x}\|_{2}\right)+(1-\gamma)\left(\|\bm{x}^{*}\|_{2}^{2}-\|\bm{x}\|_{2}^{2}\right)+\frac{1}{2\lambda}\left(\|\bm{x}^{*}-\bm{y}\|_{2}^{2}-\|\bm{x}-\bm{y}\|_{2}^{2}\right)$. Substitute the expression for $\bm{p}$ from (27) into the first term, one has $\gamma\langle\bm{p},\bm{x}^{*}-\bm{x}\rangle=\gamma\langle\frac{\bm{y}-\bm{x}^{*}}{\gamma\lambda}-\frac{2(1-\gamma)\bm{x}^{*}}{\gamma}+\frac{\alpha\bm{x}^{*}}{\|\bm{x}^{*}\|_{2}},\bm{x}^{*}-\bm{x}\rangle.$ It follows from $\frac{1}{2\lambda}\left(\|\bm{x}^{*}-\bm{y}\|_{2}^{2}-\|\bm{x}-\bm{y}\|_{2}^{2}\right)=-\frac{1}{2\lambda}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}+\frac{1}{\lambda}\langle\bm{x}^{*}-\bm{y},\bm{x}^{*}-\bm{x}\rangle$ that $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})=\gamma\frac{\alpha}{\|\bm{x}^{*}\|_{2}}\left(-\langle\bm{x}^{*},\bm{x}\rangle+\|\bm{x}^{*}\|_{2}\|\bm{x}\|_{2}\right)-(1-\gamma)\left(\|\bm{x}^{*}\|_{2}^{2}+\|\bm{x}\|_{2}^{2}\right)-\frac{1}{2\lambda}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$. From the basic inequalities $\|\bm{x}^{*}\|_{2}^{2}+\|\bm{x}\|_{2}^{2}\geq\frac{1}{2}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$ and $\|\bm{x}^{*}\|_{2}\|\bm{x}\|_{2}\leq\frac{1}{2}\|\bm{x}^{*}\|_{2}^{2}+\frac{1}{2}\|\bm{x}\|_{2}^{2}$, we have $\gamma\frac{\alpha}{\|\bm{x}^{*}\|_{2}}\left(-\langle\bm{x}^{*},\bm{x}\rangle+\|\bm{x}^{*}\|_{2}\|\bm{x}\|_{2}\right)-(1-\gamma)\left(\|\bm{x}^{*}\|_{2}^{2}+\|\bm{x}\|_{2}^{2}\right)-\frac{1}{2\lambda}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}=\frac{1}{2}\left(\frac{\gamma\alpha}{\|\bm{x}^{*}\|_{2}}-\frac{1}{\lambda}-(1-\gamma)\right)$.
@@
Case 2: $\bm{x}^{*}=\bm{0}$. It follows from $f_{\gamma,\alpha}(\bm{0})\leq f_{\gamma,\alpha}(\bm{x})$ that $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq 0$. In the special case $\alpha=0$, expanding $f_{\gamma,\alpha}(\bm{0})-f_{\gamma,\alpha}(\bm{x})$ yields $f_{\gamma,\alpha}(\bm{0})-f_{\gamma,\alpha}(\bm{x})=\frac{1}{\lambda}\langle\bm{x},\bm{y}\rangle-\gamma\|\bm{x}\|_{1}-\left(1-\gamma+\frac{1}{2\lambda}\right)\|\bm{x}\|_{2}^{2}$. According to statement (d) in Lemma 3.3, we have $\|\bm{y}\|_{\infty}\leq\gamma\lambda$, which implies that $\langle\bm{x},\bm{y}\rangle\leq\gamma\lambda\|\bm{x}\|_{1}$ and $\frac{1}{\lambda}\langle\bm{x},\bm{y}\rangle-\gamma\|\bm{x}\|_{1}\leq 0$. It follows that $f_{\gamma,\alpha}(\bm{0})-f_{\gamma,\alpha}(\bm{x})\leq-\left(1-\gamma+\frac{1}{2\lambda}\right)\|\bm{x}\|_{2}^{2}\leq 0$. Combining both cases, since $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq 0$ always holds, and in Case 1 it is bounded by a quadratic term, the uniform bound is $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq\min\left(\frac{1}{2}\left(\frac{\gamma\alpha}{\|\bm{x}^{*}\|_{2}}-\frac{1}{\lambda}-(1-\gamma)\right),\;0\right)\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$. When $\gamma=1$, Lemma 3.6 degenerates into Lemma 2 in [lou2018fast].
4 DC algorithm
@@
4.1 DC decomposition and iteration framework
@@
(28) $\mathcal{H}(\bm{x})=\underbrace{\frac{1}{2}\|\bm{Ax}-\bm{b}\|_{2}^{2}+\lambda\left(\gamma\|\bm{x}\|_{1}+\frac{3(1-\gamma)}{2}\|\bm{x}\|_{2}^{2}\right)}_{g(\bm{x})}-\underbrace{\lambda\left(\alpha\gamma\|\bm{x}\|_{2}+\frac{1-\gamma}{2}\|\bm{x}\|_{2}^{2}\right)}_{h(\bm{x})},$
which is a DC function, amenable to optimization via DCA. Let $\bm{d}_{\ell_{2}}(\bm{x}_{k})=\bigl(\alpha\gamma/\|\bm{x}_{k}\|_{2}\allowbreak+1-\gamma\bigr)\bm{x}_{k}$. The update of DCA for minimizing $\mathcal{H}(\bm{x})$ takes the form
(29) $\bm{x}_{k+1}=\begin{cases}\arg\min_{\bm{x}\in\mathbb{R}^{n}}g(\bm{x}),&\text{if }\bm{x}_{k}=0,\\ \arg\min_{\bm{x}\in\mathbb{R}^{n}}g(\bm{x})-\lambda\left\langle\bm{x},\bm{d}_{\ell_{2}}(\bm{x}_{k})\right\rangle,&\text{otherwise},\end{cases}$
@@
4.2 ADMM solver for the subproblem in DCEN-DCA

### v2

@@ IV Proximal Operator
$\operatorname{prox}_{\lambda r_{\gamma}}(\bm{y})=\arg\min_{\bm{x}\in\mathbb{R}^{n}}f_{\gamma,\alpha}(\bm{x})=r(\bm{x})+\frac{1}{2\lambda}\|\bm{x}-\bm{y}\|_{2}^{2},$ (31)
@@
Theorem 11 (Solution’s Existence).
@@
Proof.
Clearly, the objective function in (31) is proper and closed. It follows from the inequality $\|\bm{x}\|_{1}\geq\|\bm{x}\|_{2}$ that $f_{\gamma,\alpha}(\bm{x})\geq\left(1-\gamma+\frac{1}{2\lambda}\right)\|\bm{x}\|_{2}^{2}+\left[\gamma(1-\alpha)-\frac{1}{\lambda}\|\bm{y}\|_{2}\right]\|\bm{x}\|_{2}+\frac{1}{2\lambda}\|\bm{y}\|_{2}^{2}$. The right-hand side is a quadratic function of $\|\bm{x}\|_{2}$ with a positive leading coefficient $1-\gamma+\frac{1}{2\lambda}>0$. Hence, $f_{\gamma,\alpha}(\bm{x})\to+\infty$ as $\|\bm{x}\|_{2}\to\infty$, which implies that $f_{\gamma,\alpha}$ is coercive. Since $f_{\gamma,\alpha}$ is also lower semicontinuous, its sublevel sets are compact and so $\operatorname{prox}_{\lambda r_{\gamma}}(\bm{y})\neq\emptyset$. ∎
Lemma 12.
@@
    If $\|\bm{y}\|_{\infty}>\gamma\lambda$, then $\bm{x}^{*}=\frac{\|\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})\|_{2}+\alpha\gamma\lambda}{(1+2\lambda(1-\gamma))\,\|\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})\|_{2}}\;\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})$, where $\mathcal{\bm{S}}_{\gamma\lambda}(\bm{y})=\operatorname{sign}(\bm{y})\odot\max\big(|\bm{y}|-\gamma\lambda,\,\bm{0}\big)$.
@@
    If $\|\bm{y}\|_{\infty}=\gamma\lambda$, then the optimal solution $\bm{x}^{*}$ satisfies $\bm{x}^{*}\in\{\bm{x}\in\mathbb{R}^{n}:x_{i}=0\text{ for }|y_{i}|<\gamma\lambda,\ \|\bm{x}\|_{2}=\alpha\gamma\lambda,\ x_{i}y_{i}\geq 0,\forall i\}$, and $\bm{x}^{*}$ is non-unique if multiple $|y_{i}|=\gamma\lambda$.
@@
    If $(1-\alpha)\gamma\lambda<\|\bm{y}\|_{\infty}<\gamma\lambda$, then $\bm{x}^{*}$ is $1$-sparse, supported on an index $i$ with $|y_{i}|=\|\bm{y}\|_{\infty}$, and $\bm{x}^{*}\in\{\bm{x}\in\mathbb{R}^{n}\Big|\|\bm{x}\|_{2}=\|\bm{y}\|_{\infty}+(\alpha-1)\gamma\lambda,\ x_{i}y_{i}\geq 0\}.$
@@
Proof.
By combining the quadratic terms in the proximal mapping (31), the resulting operator can be expressed in the same form as the $\ell_{1}-\alpha\ell_{2}$ proximal mapping $\operatorname{prox}_{\lambda r_{\gamma}}(\bm{\tilde{y}})=\arg\min_{\bm{x}\in\mathbb{R}^{n}}\,\|\bm{x}\|_{1}-\alpha\|\bm{x}\|_{2}+\frac{1}{2\tilde{\lambda}}\|\bm{x}-\bm{\tilde{y}}\|_{2}^{2}$, where $\tilde{\lambda}=\frac{\lambda\gamma}{1+2\beta(1-\gamma)}$ and $\bm{\tilde{y}}=\frac{1}{1+2\beta(1-\gamma)}\bm{y}$. This reformulation allows one to directly apply the closed-form expressions derived for the $\ell_{1}-\alpha\ell_{2}$ case. The remaining proof follows that of [25] and is therefore omitted. ∎
Lemma 13.
@@
$\bm{p}=\frac{1}{\lambda\gamma}\bm{y}-\left(\frac{1}{\lambda\gamma}+\frac{2(1-\gamma)}{\gamma}-\frac{\gamma\alpha}{\|\bm{x}^{*}\|_{2}}\right)\bm{x}^{*}.$ (32)
@@
By the subgradient inequality for $\|\cdot\|_{1}$, one has $\|\bm{x}\|_{1}\geq\|\bm{x}^{*}\|_{1}+\langle\bm{p},\bm{x}-\bm{x}^{*}\rangle$, which implies $\|\bm{x}^{*}\|_{1}-\|\bm{x}\|_{1}\leq\langle\bm{p},\bm{x}^{*}-\bm{x}\rangle$ and so $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq\gamma\langle\bm{p},\bm{x}^{*}-\bm{x}\rangle-\gamma\alpha\left(\|\bm{x}^{*}\|_{2}-\|\bm{x}\|_{2}\right)+(1-\gamma)\left(\|\bm{x}^{*}\|_{2}^{2}-\|\bm{x}\|_{2}^{2}\right)+\frac{1}{2\lambda}\left(\|\bm{x}^{*}-\bm{y}\|_{2}^{2}-\|\bm{x}-\bm{y}\|_{2}^{2}\right)$. Substitute the expression for $\bm{p}$ from (32) into the first term, one has $\gamma\langle\bm{p},\bm{x}^{*}-\bm{x}\rangle=\gamma\langle\frac{\bm{y}-\bm{x}^{*}}{\gamma\lambda}-\frac{2(1-\gamma)\bm{x}^{*}}{\gamma}+\frac{\alpha\bm{x}^{*}}{\|\bm{x}^{*}\|_{2}},\bm{x}^{*}-\bm{x}\rangle.$ It follows from $\frac{1}{2\lambda}\left(\|\bm{x}^{*}-\bm{y}\|_{2}^{2}-\|\bm{x}-\bm{y}\|_{2}^{2}\right)=-\frac{1}{2\lambda}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}+\frac{1}{\lambda}\langle\bm{x}^{*}-\bm{y},\bm{x}^{*}-\bm{x}\rangle$ that $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})=\gamma\frac{\alpha}{\|\bm{x}^{*}\|_{2}}\left(-\langle\bm{x}^{*},\bm{x}\rangle+\|\bm{x}^{*}\|_{2}\|\bm{x}\|_{2}\right)-(1-\gamma)\left(\|\bm{x}^{*}\|_{2}^{2}+\|\bm{x}\|_{2}^{2}\right)-\frac{1}{2\lambda}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$. From the basic inequalities $\|\bm{x}^{*}\|_{2}^{2}+\|\bm{x}\|_{2}^{2}\geq\frac{1}{2}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$ and $\|\bm{x}^{*}\|_{2}\|\bm{x}\|_{2}\leq\frac{1}{2}\|\bm{x}^{*}\|_{2}^{2}+\frac{1}{2}\|\bm{x}\|_{2}^{2}$, we have
$\gamma\frac{\alpha}{\|\bm{x}^{*}\|_{2}}\left(-\langle\bm{x}^{*},\bm{x}\rangle+\|\bm{x}^{*}\|_{2}\|\bm{x}\|_{2}\right)$
$\quad-(1-\gamma)\left(\|\bm{x}^{*}\|_{2}^{2}+\|\bm{x}\|_{2}^{2}\right)-\frac{1}{2\lambda}\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$
$\quad\quad=\frac{1}{2}\left(\frac{\gamma\alpha}{\|\bm{x}^{*}\|_{2}}-\frac{1}{\lambda}-(1-\gamma)\right).$
@@
Case 2: $\bm{x}^{*}=\bm{0}$. It follows from $f_{\gamma,\alpha}(\bm{0})\leq f_{\gamma,\alpha}(\bm{x})$ that $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq 0$. In the special case $\alpha=0$, expanding $f_{\gamma,\alpha}(\bm{0})-f_{\gamma,\alpha}(\bm{x})$ yields $f_{\gamma,\alpha}(\bm{0})-f_{\gamma,\alpha}(\bm{x})=\frac{1}{\lambda}\langle\bm{x},\bm{y}\rangle-\gamma\|\bm{x}\|_{1}-\left(1-\gamma+\frac{1}{2\lambda}\right)\|\bm{x}\|_{2}^{2}$. According to statement (d) in Lemma 12, we have $\|\bm{y}\|_{\infty}\leq\gamma\lambda$, which implies that $\langle\bm{x},\bm{y}\rangle\leq\gamma\lambda\|\bm{x}\|_{1}$ and $\frac{1}{\lambda}\langle\bm{x},\bm{y}\rangle-\gamma\|\bm{x}\|_{1}\leq 0$. It follows that $f_{\gamma,\alpha}(\bm{0})-f_{\gamma,\alpha}(\bm{x})\leq-\left(1-\gamma+\frac{1}{2\lambda}\right)\|\bm{x}\|_{2}^{2}\leq 0$. Combining both cases, since $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq 0$ always holds, and in Case 1 it is bounded by a quadratic term, the uniform bound is $f_{\gamma,\alpha}(\bm{x}^{*})-f_{\gamma,\alpha}(\bm{x})\leq\min\left(\frac{1}{2}\left(\frac{\gamma\alpha}{\|\bm{x}^{*}\|_{2}}-\frac{1}{\lambda}-(1-\gamma)\right),\;0\right)\|\bm{x}^{*}-\bm{x}\|_{2}^{2}$. When $\gamma=1$, Lemma 13 degenerates into Lemma 2 in [25]. ∎
V DC algorithm
@@
V-A DC decomposition and iteration framework
@@
$\mathcal{H}(\bm{x})=g(\bm{x})-h(\bm{x}),$ (33)
where $g(\bm{x})=\ell(\bm{x})+\lambda(\gamma\|\bm{x}\|_{1}+\frac{3(1-\gamma)}{2}\|\bm{x}\|_{2}^{2})$ and $h(\bm{x})=\lambda(\alpha\gamma\|\bm{x}\|_{2}+\frac{1-\gamma}{2}\|\bm{x}\|_{2}^{2})$. (33) is a DC function, amenable to optimization via DCA. Let $\bm{d}_{\ell_{2}}(\bm{x}_{k})=\bigl(\alpha\gamma/\|\bm{x}_{k}\|_{2}\allowbreak+1-\gamma\bigr)\bm{x}_{k}$. The update of DCA for minimizing $\mathcal{H}(\bm{x})$ takes the form
$\bm{x}_{k+1}=\begin{cases}\arg\min_{\bm{x}\in\mathbb{R}^{n}}g(\bm{x}),&\text{if }\bm{x}_{k}=0,\\ \arg\min_{\bm{x}\in\mathbb{R}^{n}}g(\bm{x})-\lambda\left\langle\bm{x},\bm{d}_{\ell_{2}}(\bm{x}_{k})\right\rangle,&\text{otherwise},\end{cases}$ (34)
@@
V-B ADMM solver for the subproblem in DCEN-DCA